# The Onset of Intermittency in Active Turbulence

### v1

@@ Abstract
Figure 2: Probability density functions (PDFs) of the Lagrangian velocity increments $\Delta v_{i}$ (both Cartesian coordinates included) for different time lags $\tau$ (see legend) and for $\alpha=-6$. Clearly, these distributions are Gaussian (indicated by the dashed line) at large time lags, but show significant non-Gaussian tails as $\tau$ becomes smaller. (The curves are staggered vertically for clarity.)
@@
Before we turn to the actual measurements of the Lagrangian exponents, we recall that there is a simple dimensional argument Biferale et al. (2004); Arnèodo et al. (2008) which leads to a relation between the Lagrangian exponents $\xi_{p}$ and the Eulerian exponents $\zeta_{p}$: If in time $\tau$, the Lagrangian particle moves from ${\bf x}_{1}$ to ${\bf x}_{2}$, with ${\bf r}=|{\bf x}_{2}-{\bf x}_{1}|$ and $r=|{\bf r}|$, and we assume that the corresponding Lagrangian-velocity increment $\Delta v\sim|\Delta{\bf u}\cdot\hat{\bf r}|\sim S_{1}(r)$, then it follows that $\Delta v\sim r^{\zeta_{1}}$. The estimate $r\sim\Delta v/\tau$ for the distance traveled by the Lagrangian particle , yields $\Delta v\sim\tau^{\frac{\zeta_{1}}{1-\zeta_{1}}}$; this implies that $\xi_{1}=\frac{\zeta_{1}}{1-\zeta_{1}}$. If we assume that simple scaling holds (i.e., there is no intermittency), then $\zeta_{p}=p/4$; and $\xi_{1}=1/3$ and $\xi_{p}=p/3$. (For inertial turbulence, a similar argument gives $\xi_{p}=p/2$ Biferale et al. (2004); Arnèodo et al. (2008).)
@@
We first focus on the second-order Lagrangian structure function to confirm $\xi_{2}\simeq 2/3$. Then we use a combination of ESS Benzi et al. (1993); Chakraborty et al. (2010) and local-slope analysis to calculate $\xi_{p}/\xi_{2}$. In the inset of Fig. 3, we show (for $\alpha=-6$) plots of the local slopes $\frac{d\log S^{\rm L}_{p}}{d\log S^{\rm L}_{2}}$ vs. $\tau$. In the limit $\tau\to 0$, the local slopes asymptote $\sim p/2$ beacuse small-scale smoothness yields the Taylor-expansion result $S^{\rm L}_{p}\sim\tau^{p}$. For slightly larger values of $\tau\simeq\mathcal{O}(10^{-1})$, there is a dip in the local slopes, which is familiar from similar measurements in inertial turbulence Arnèodo et al. (2008); Buzzicotti et al. (2016)), where this dip is attributed to the trapping of tracers in vortices. For even larger values of $\tau$, beyond this dip, a new scaling range emerges for $S^{\rm L}_{p}(\tau)$, which corresponds to the scaling range for the Eulerian $S_{p}(r)$ [Fig. 1], if we relate time and length scales via $\tau\sim\Delta v/r$. In the inset Fig. 3, this range seen appears as a plateau after the dip; the mean value of this plateau and the standard deviation from this mean yield the exponent ratios $\xi_{p}/\xi_{2}$ and their error bars, respectively. If there were no intermittency, we would have found $\xi_{p}=p/3$ and $\xi_{p}/\xi_{2}=p/2$. Our plot the exponent ratios as a function of $p$, in Fig. 3, shows significant deviations from the simple-scaling the $p/2$ line. This clear departure from linear scaling underlines the presence of Lagrangian intermittency in low-Re TTSH turbulence. Furthermore, we find another intriguing similarity with recent results for inertial turbulence Iyer et al. (2020); Buaria and Sreenivasan (2023b): The exponents $\xi_{p}$ saturate to a constant for $p\gtrsim 5$. In inertial turbulence, this saturation of Lagrangian exponents has been attributed to the saturation of the exponents $\zeta^{\perp}_{p}$ of the transverse, equal-time Eulerian structure functions: $S^{\perp}_{p}(r)\equiv\langle|\Delta{\bf u}-(\Delta{\bf u}\cdot\hat{\bf r})\hat{\bf r}|^{p}\rangle\sim r^{\zeta^{\perp}_{p}}$. We uncover a similar correspondence with $\zeta^{\perp}_{p}\approx\xi_{p}$ for $p\gtrsim 5$ (see Fig. 3); these exponent ratios appear to saturate to $\simeq 1.5$. In inertial turbulence, it has been argued such saturation is likely because of slender vortex filaments Frisch (1996); Buaria and Sreenivasan (2023b). We conjecture that the exponent saturation we see, for $\alpha\lesssim\alpha_{c}$, may be related to localised streak-like structures, which have been seen in earlier studies Mukherjee et al. (2021); Singh et al. (2022); Puggioni et al. (2022).
@@ Acknowledgements.
RP and KVK acknowledge support from SERB and NSM (India) and thank SERC (IISc) for computational resources. KK and SSR are indebted to S. Mukherjee for his tutelage on understanding active turbulence as well as developing the DNS code at ICTS. The CPU/MPI simulations were performed on the ICTS clusters Tetris and Contra. The GPU simulations were performed on the CUDA A100 at Departement of Physics, IISc. SSR acknowledges SERB-DST (India) projects STR/2021/000023 and CRG/2021/002766 for financial support. KK and SSR acknowledges the support of the DAE, Govt. of India, under project no. 12-R&D-TFR-5.10-1100 and project no. RTI4001. AG acknowledges SERB-DST (India) projects MTR/2022/000232, CRG/2023/007056-G for financial support.

### v2

@@ Abstract
Figure 2: Probability density functions (PDFs) of the Lagrangian velocity increments $\Delta v_{i}$ (both Cartesian coordinates included) for different time lags $\tilde{\tau}=\tau/(2\sqrt{2}\pi\Gamma_{2}/|\Gamma_{0}|^{2})$ (see legend) and for $\alpha=-6$. Clearly, these distributions are Gaussian (indicated by the dashed line) at large time lags, but show significant non-Gaussian tails as $\tau$ becomes smaller. (The curves are staggered vertically for clarity.)
@@
Before we turn to the actual measurements of the Lagrangian exponents, we recall that there is a simple dimensional argument Biferale et al. (2004); Arnèodo et al. (2008) which leads to a relation between the Lagrangian exponents $\xi_{p}$ and the Eulerian exponents $\zeta_{p}$: If in time $\tau$, the Lagrangian particle moves from ${\bf x}_{1}$ to ${\bf x}_{2}$, with ${\bf r}=|{\bf x}_{2}-{\bf x}_{1}|$ and $r=|{\bf r}|$, and we assume that the corresponding Lagrangian-velocity increment $\Delta v\sim|\Delta{\bf u}\cdot\hat{\bf r}|\sim S_{1}(r)$, then it follows that $\Delta v\sim r^{\zeta_{1}}$. The estimate $r\sim(\Delta v)\tau$ for the distance traveled by the Lagrangian particle , yields $\Delta v\sim\tau^{\frac{\zeta_{1}}{1-\zeta_{1}}}$; this implies that $\xi_{1}=\frac{\zeta_{1}}{1-\zeta_{1}}$. If we assume that simple scaling holds (i.e., there is no intermittency), then $\zeta_{p}=p/4$; and $\xi_{1}=1/3$ and $\xi_{p}=p/3$. (For inertial turbulence, a similar argument gives $\xi_{p}=p/2$ Biferale et al. (2004); Arnèodo et al. (2008).)
@@
We first focus on the second-order Lagrangian structure function to confirm $\xi_{2}\simeq 2/3$. Then we use a combination of ESS Benzi et al. (1993); Chakraborty et al. (2010) and local-slope analysis to calculate $\xi_{p}/\xi_{2}$. In the inset of Fig. 3, we show (for $\alpha=-6$) plots of the local slopes $\frac{d\log S^{\rm L}_{p}}{d\log S^{\rm L}_{2}}$ vs. $\tau$. In the limit $\tau\to 0$, the local slopes asymptote $\sim p/2$ beacuse small-scale smoothness yields the Taylor-expansion result $S^{\rm L}_{p}\sim\tau^{p}$. For slightly larger values of $\tau\simeq\mathcal{O}(10^{-1})$, there is a dip in the local slopes, which is familiar from similar measurements in inertial turbulence Arnèodo et al. (2008); Buzzicotti et al. (2016)), where this dip is attributed to the trapping of tracers in vortices. For even larger values of $\tau$, beyond this dip, a new scaling range emerges for $S^{\rm L}_{p}(\tau)$, which corresponds to the scaling range for the Eulerian $S_{p}(r)$ [Fig. 1], if we relate time and length scales via $\tau\sim r/\Delta v$. In the inset Fig. 3, this range seen appears as a plateau after the dip; the mean value of this plateau and the standard deviation from this mean yield the exponent ratios $\xi_{p}/\xi_{2}$ and their error bars, respectively. If there were no intermittency, we would have found $\xi_{p}=p/3$ and $\xi_{p}/\xi_{2}=p/2$. Our plot the exponent ratios as a function of $p$, in Fig. 3, shows significant deviations from the simple-scaling the $p/2$ line. This clear departure from linear scaling underlines the presence of Lagrangian intermittency in low-Re TTSH turbulence. Furthermore, we find another intriguing similarity with recent results for inertial turbulence Iyer et al. (2020); Buaria and Sreenivasan (2023b): The exponents $\xi_{p}$ saturate to a constant for $p\gtrsim 5$. In inertial turbulence, this saturation of Lagrangian exponents has been attributed to the saturation of the exponents $\zeta^{\perp}_{p}$ of the transverse, equal-time Eulerian structure functions: $S^{\perp}_{p}(r)\equiv\langle|\Delta{\bf u}-(\Delta{\bf u}\cdot\hat{\bf r})\hat{\bf r}|^{p}\rangle\sim r^{\zeta^{\perp}_{p}}$. We uncover a similar correspondence with $\zeta^{\perp}_{p}\approx\xi_{p}$ for $p\gtrsim 5$ (see Fig. 3); these exponent ratios appear to saturate to $\simeq 1.5$. In inertial turbulence, it has been argued such saturation is likely because of slender vortex filaments Frisch (1996); Buaria and Sreenivasan (2023b). We conjecture that the exponent saturation we see, for $\alpha\lesssim\alpha_{c}$, may be related to localised streak-like structures, which have been seen in earlier studies Mukherjee et al. (2021); Singh et al. (2022); Puggioni et al. (2022).
@@ Acknowledgements.
RP and KVK acknowledge support from SERB and NSM (India) and thank SERC (IISc) for computational resources. KK and SSR are indebted to S. Mukherjee for his tutelage on understanding active turbulence as well as developing the DNS code at ICTS. The CPU/MPI simulations were performed on the ICTS clusters Tetris and Contra. The GPU simulations were performed on the CUDA A100 at Departement of Physics, IISc. KVK and SSR thank Rahul Singh for careful reading of the manuscript. SSR acknowledges SERB-DST (India) projects STR/2021/000023 and CRG/2021/002766 for financial support. KK and SSR acknowledges the support of the DAE, Govt. of India, under project no. 12-R&D-TFR-5.10-1100 and project no. RTI4001. AG acknowledges SERB-DST (India) projects MTR/2022/000232, CRG/2023/007056-G for financial support.